# What Does Stable Diffusion Know about the 3D Scene?

### v2

@@ Abstract
Recent advances in generative models like Stable Diffusion (Rombach et al., 2022) enable the generation of highly photo-realistic images. Our objective in this paper is to probe the diffusion network to determine to what extent it ‘understands’ different properties of the 3D scene depicted in an image. To this end, we make the following contributions: (i) We introduce a protocol to evaluate whether features of an off-the-shelf diffusion model encode a number of physical ‘properties’ of the 3D scene, by training discriminative classifiers on the features for these properties. The probes are applied on datasets of real images with annotations for the property. (ii) We apply this protocol to properties covering scene geometry, scene material, support relations, lighting, and view dependent measures. (iii) We find that features from Stable Diffusion are good for discriminative learning of a number of properties, including scene geometry, support relations, shadows and depth, but less performant for occlusion and material. (iv) We also apply the probes to other networks trained at large-scale, including DINO, CLIP and VQGAN, and find that DINOv2 has a similar performance to Stable Diffusion, while outperforming DINOv1, CLIP and VQGAN.
3D Physical Scene Understanding, Stable Diffusion, Representation Learning
@@ 1 Introduction
Figure 1: Motivation: What does Stable Diffusion know about the 3D scene? The Stable Diffusion inpainting model is here tasked with inpainting the masked region of the real images. It correctly predicts a shadow consistent with the lighting direction (top), and a supporting structure consistent with the scene geometry (bottom). This indicates that the Stable Diffusion model generation is consistent with the geometry (of the light source direction) and physical (support) properties. Note: these examples are only for illustration and inpainting is not the objective of this paper. Rather, we probe a general Stable Diffusion network to determine whether there are explicit features for such 3D scene properties. The appendix provides more examples of Stable Diffusion’s capability to predict different physical properties of the scene.
@@
Image generation with diffusion models (Sohl-Dickstein et al., 2015), following on from earlier generation using GANs (Goodfellow et al., 2014), has achieved amazing results in terms of verisimilitude (Rombach et al., 2022). This naturally raises the question of to what extent the diffusion network is modelling the 3D scene depicted in the image. Our objective in this paper is to investigate this question, and we do this precisely by determining whether features from the diffusion network can be used to predict the true geometric and physical properties of the 3D scene. If they can, then that is evidence that the 3D scene is correctly modelled. As an indication that the diffusion network is 3D and physics aware, Figure 1 shows the result of the off-the-shelf Stable Diffusion model (Rombach et al., 2022) inpainting masked regions in real images – it correctly predicts shadows and supporting structures.
@@
To answer this question, we propose an evaluation protocol to systematically ‘probe’ a diffusion network on its ability to represent a number of ‘properties’ of the 3D scene and viewpoint. These properties include: 3D structure and material of the scene, such as surface layout; lighting, such as object-shadow relationships; and viewpoint dependent relations such as occlusion and depth.
@@
The protocol involves three steps: First, a suitable real image evaluation dataset is selected that contains ground truth annotations for the property of interest, for example the SOBA dataset (Wang et al., 2020) is used to probe the understanding of lighting, as it has annotations for object-shadow associations. This provides a train/val/test set for that property; Second, a grid search is carried out over the layers and time steps of the diffusion model to select the optimal feature for determining that property. The selection involves learning the weights of a simple linear classifier for that property (e.g. ‘are these two regions in an object-shadow relationship or not’); Third, the selected feature (layer, time step) and trained classifier are evaluated on a test set, and its performance answers the question of how well the diffusion model ‘understands’ that property.
@@
In short, we probe scene geometry, material, support relation, shadow, occlusion and depth, to answer the question “What does Stable Diffusion Know about the 3D scene?”. We apply this protocol not only to features from the Stable Diffusion network, but also to features from other networks that are trained at large scale, including OpenCLIP (Radford et al., 2021; Ilharco et al., 2021), DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and VQGAN (Esser et al., 2021). This covers networks trained generatively (Stable Diffusion), with self-supervision (DINOv1 & DINOv2), with weak supervision (OpenCLIP), and by auto-regression (VQGAN).
@@
From our investigation, we make two observations: First, the Stable Diffusion and DINOv2 networks have a good ‘understanding’ of the scene geometry, support relations, the lighting, and the depth of a scene, with Stable Diffusion and DINOv2 having a similar and high prediction performance for these properties. However, their prediction of material and occlusion is poorer. Second, Stable Diffusion and DINOv2 generally demonstrate better performance for 3D properties, than other networks trained at large scale: OpenCLIP, DINOv1, and VQGAN.
@@
Why is an understanding of the networks’ ability to predict 3D properties useful? There are four reasons: (1) It begins to answer the scientific question of the extent to which these networks implicitly model the 3D scene; (2) The features we determine that are able to predict 3D physical properties can be used for this task, e.g. to predict shadow-object associations or support relations. This could either be carried out directly by incorporating them in a prediction network, in the manner of Zhan et al. (2023); or they can be used indirectly as a means to train a feed forward network to predict the properties (Wu et al., 2023a, b); (3) By knowing what properties Stable Diffusion is not good at, we have a way to spot images generated by Stable Diffusion, as has been done by Sarkar et al. (2023); (4) It also reveals which properties the network could be trained further on to improve its 3D modelling, e.g., via extra supervision for that property.
We describe the properties explored, the protocol, datasets and classifiers in Section 3. Experimental results of both Stable Diffusion and other networks are given in Section 4, and finally future work is discussed in Section 5.
@@
2 Related Work
2.1 Generative Models
Generative models have made significant achievements in advancing image quality and diversity in the recent literature. A series of generative models, such as Variational Autoencoders (VAEs) (Kingma & Welling, 2014), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Flow-based Generators (Dinh et al., 2014), and Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015), have been proposed. These models have contributed to widespread tasks, including image completion (Pathak et al., 2016), composition (Lin et al., 2018), interpolation (Karras et al., 2019) and editing (Chai et al., 2021), image-to-image translation (Isola et al., 2017), multi-modalities translation (Hu et al., 2023), and numerous others. We build upon the diffusion models (Rombach et al., 2022), which have shown state-of-the-art generation quality.
2.2 Exploration of Pre-trained Models
Building on the success of deep networks, there has been significant interest from the community to understand what has been learnt by these complex models. On discriminative models, for example, Zeiler & Fergus (2014); Mahendran & Vedaldi (2015) propose inverse reconstruction to directly visualize the acquired semantic information in various layers of a trained classification network; Zhou et al. (2016); Fong & Vedaldi (2017); Fong et al. (2019) demonstrate that scene classification networks have remarkable localization ability despite being trained on only image-level labels; and Erhan et al. (2009); Simonyan et al. (2014); Selvaraju et al. (2017) use the gradients of any target concept, flowing into the final convolutional layer to produce a saliency map highlighting important regions in the image for predicting the concept. In the more recent literature, Chefer et al. (2021) explore what has been learned in the powerful transformer model by visualizing the attention map. On generative models, researchers have mainly investigated what has been learned in GANs, for example, GAN dissection (Bau et al., 2019) presents an analytic framework to visualize and understand GANs at the unit-, object-, and scene-level; Wu et al. (2021) analyse the latent style space of StyleGANs (Karras et al., 2019). The most recent work (Sarkar et al., 2023) studies the 3D geometric relations in generated images, such as vanishing points and shadows, and notes that the errors made can be used to discriminate real from generated images.
2.3 Exploitation of Generative Models
Apart from understanding the representation in pre-trained models, there has been a recent trend for exploiting the learnt feature from generative models, to tackle a series of downstream discriminative tasks. For example, leveraging generative models for data augmentation in recognition tasks (Jahanian et al., 2022; He et al., 2023), semantic segmentation via generative models (Li et al., 2021; Baranchuk et al., 2022; Xu et al., 2023), open-vocabulary segmentation with diffusion models (Li et al., 2023), depth maps estimation based on RGB images (Noguchi & Harada, 2020; Shi et al., 2022). More recently, Bhattad et al. (2023) search for intrinsic offsets in a pre-trained StyleGAN for a range of downstream tasks, predicting normal maps, depth maps, segmentations, albedo maps, and shading. In contrast to this work, we adopt annotations from different datasets for supervision, rather than employing pre-trained prediction models for supervision. A closely related effort to ours is DIffusion FeaTures (DIFT) (Tang et al., 2023), but it only focuses on computing correspondences at the geometric or semantic level between images.
2.4 Physical Scene Understanding
There have been works studying different physical properties for scene understanding, including shadows (Wang et al., 2020, 2021), material (Upchurch & Niu, 2022; Sharma et al., 2023), occlusion (Zhan et al., 2022), scene geometry (Liu et al., 2019), support relations (Silberman et al., 2012) and depth (Silberman et al., 2012). However, these works focus on one or two physical properties, and most of them require training a model for the property in a supervised manner. In contrast, we use a single model to predict multiple properties, and do not train the features.
3 Method – Properties, Datasets, and Classifiers
Figure 2: Example images for probing scene geometry. The first row shows a sample annotation for the same plane, and the second row is a sample annotation for perpendicular plane. Here, and in the following figures, (A, B) are a positive pair, while (A, C) are negative. The images are from the ScanNetv2 dataset (Dai et al., 2017) with annotations for planes from (Liu et al., 2019). In the first row, the first piece of floor (A) is on the same plane as the second piece of floor (B), but is not on the same plane as the surface of the drawers (C). In the second row, the table top (A) is perpendicular to the wall (B), but is not perpendicular to the stool top (C).
Figure 3: Example images for probing material, support relation and shadow. The first row is for material, the second row for support relation, and the third row for shadow. First row: the material images are from the DMS dataset (Upchurch & Niu, 2022). The paintings are both covered with glass (A and B) whereas the curtain (C) is made of fabric. Second row: the support relation images are from the NYUv2 dataset (Silberman et al., 2012). The paper (A) is supported by the table (B), but it is not supported by the chair (C). Third row: the shadow images are from the SOBA dataset (Wang et al., 2020). The person (A) has the shadow (B), not the shadow (C).
Our goal is to examine the ability of a diffusion model to understand different physical properties of the 3D scene, including: scene geometry, material, support relations, shadows, occlusion and depth. Specifically, we conduct linear probing of the features from different layers and time steps of the Stable Diffusion model. First, we set up the questions for each property (Section 3.1); and then select real image datasets with ground truth annotations for each property (Section 3.2). We describe how a classifier is trained to answer the questions, and the grid search for the optimal time step and layer to extract a feature for predicting the property in Section 3.3.
3.1 Properties and Questions
Here, we study the diffusion model’s ability to predict different properties of the 3D scene; the properties cover the 3D structure and material, the lighting, and the viewpoint. For each property, we propose questions that classify the relationship between a pair of Regions, A and B, in the same image, based on the features extracted from the diffusion model. The properties and questions are:
  • Same Plane: ‘Are Region A and Region B on the same plane?’
  • Material: ‘Are Region A and Region B made of the same material?’
  • Shadow: ‘Are Region A and Region B in an object-shadow relationship?’
  • Depth: ‘Does Region A have a greater average depth than Region B?’
@@ 3.2 Datasets
Same Plane. We use the ScanNetv2 dataset (Dai et al., 2017) with annotations for planes from Liu et al. (2019). Regions are obtained via splitting plane masks into several regions. A pair of regions are positive if they are on the same plane, and negative if they are on different planes. First row of Figure 2 is an example.
@@
Perpendicular Plane. We use the ScanNetv2 dataset (Dai et al., 2017). We use the annotations from Liu et al. (2019) which provide segmentation masks as well as plane parameters for planes in the image, so that we can obtain the normal of planes to judge whether they are perpendicular or not. A pair of regions are positive if they are on perpendicular planes, and negative if they are not on perpendicular planes. Second row of Figure 2 is an example.
@@
Material. We adopt the recent DMS dataset (Upchurch & Niu, 2022) to study the material property, which provides dense annotations of material category for each pixel in the images. Therefore, we can get the mask of each material via grouping pixels with the same material label together. In total, there are 46 pre-defined material categories. Regions are obtained by splitting the mask of each material into different connected components, i.e., they are simply groups with same material labels, yet not connected. A pair of regions are positive if they are of the same material category, and negative if they are of different material categories. First row of Figure 3 is an example.
@@
Support Relation. We use the NYUv2 dataset (Silberman et al., 2012) to probe the support relation. Segmentation annotations for different regions (objects or surfaces) are provided, as well as their support relations. Support relation here means an object is physically supported by another object, i.e., the second object will undertake the force to enable the first object to stay at its position. Regions are directly obtained via the segmentation annotations. A pair of regions are positive if the first region is supported by the second region, and negative if the first region is not supported by the second region. Second row of Figure 3 is an example.
@@
Shadow. We use the SOBA dataset (Wang et al., 2020, 2021) to study the shadows which depend on the lighting of the scene. Segmentation masks for each object and shadow, as well as their associations are provided in the dataset annotations. Regions are directly obtained from the annotated object and shadow masks. In a region pair, there is one object mask and one shadow mask. A pair of regions are positive if the shadow mask is the shadow of the object, and negative if the shadow mask is the shadow of another object. Third row of Figure 3 is an example.
@@
Figure 4: Example images for probing viewpoint-dependent properties (occlusion & depth). The first row is for occlusion and the second row is for depth. First row: the occlusion images are from the Separated COCO dataset (Zhan et al., 2022). The sofa (A) and the sofa (B) are part of the same object, whilst the monitor (C) is not part of the sofa. Second row: the depth images are from the NYUv2 dataset (Silberman et al., 2012). The chair (A) is farther away than the object on the floor (B), but it is closer than the cupboard (C).
@@
Occlusion. We use the Seperated COCO dataset (Zhan et al., 2022) to study the occlusion (object seperation) problem. Regions are different connected components of objects (and the object mask if it is not separated), i.e., groups of connected pixels belonging to the same object. A pair of regions are positive if they are different components of the same object separated due to occlusion, and negative if they are not from the same object. First row of Figure 4 is an example.
@@
Depth. We use the NYUv2 dataset (Silberman et al., 2012), that provides mask annotations for different objects and regions, together with depth for each pixel. A pair of regions are positive if the first region has a greater average depth than the second region, and negative if the first region has a less average depth than the second region. The average depth of a region is calculated via the average of depth value of each pixel the region contains. Second row of Figure 4 is an example.

### v3

# A General Protocol to Probe Large Vision Models for 3D Physical Understanding

@@ Abstract
Our objective in this paper is to probe large vision models to determine to what extent they ‘understand’ different physical properties of the 3D scene depicted in an image. To this end, we make the following contributions: (i) We introduce a general and lightweight protocol to evaluate whether features of an off-the-shelf large vision model encode a number of physical ‘properties’ of the 3D scene, by training discriminative classifiers on the features for these properties. The probes are applied on datasets of real images with annotations for the property. (ii) We apply this protocol to properties covering scene geometry, scene material, support relations, lighting, and view-dependent measures, and large vision models including CLIP, DINOv1, DINOv2, VQGAN, Stable Diffusion. (iii) We find that features from Stable Diffusion and DINOv2 are good for discriminative learning of a number of properties, including scene geometry, support relations, shadows and depth, but less performant for occlusion and material, while outperforming DINOv1, CLIP and VQGAN for all properties. (iv) It is observed that different time steps of Stable Diffusion features, as well as different transformer layers of DINO/CLIP/VQGAN, are good at different properties, unlocking potential applications of 3D physical understanding.
@@ 1 Introduction
Figure 1: Motivation: What do large vision models know about the 3D scene? We take Stable Diffusion as an example because Stable Diffusion is generative, and so its output is an image that can be judged directly for verisimilitude. The Stable Diffusion inpainting model is here tasked with inpainting the masked region of the real images. It correctly predicts a shadow consistent with the lighting direction (top), and a supporting structure consistent with the scene geometry (bottom). This indicates that the Stable Diffusion model generation is consistent with the geometry (of the light source direction) and physical (support) properties. These examples are only for illustration and we probe a general Stable Diffusion network to determine whether there are explicit features for such 3D scene properties. The appendix provides more examples of Stable Diffusion’s capability to predict different physical properties of the scene.
@@
The large-scale pre-trained vision foundation models have achieved great success in computer vision tasks, including classification (CLIP [30, 18]), segmentation (DINO [5, 27]), and image generation (VQGAN [11], Stable Diffusion [31]) with strong generalisation capabilities. However, they are mainly trained with 2D images, which are the projection of the 3D physical world. This naturally raises the question of to what extent these large-scale vision models have learned about the 3D scene depicted with only the 2D images. Our objective in this paper is to investigate this question, and we do this precisely by determining whether features from these large-scale pre-trained vision models can be used to estimate the true geometric and physical properties of the 3D scene. If they can, then that is evidence that the 3D scene can be correctly modelled by its 2D projections. For example, as an indication that Stable Diffusion is 3D and physics aware, Figure 1 shows the result of the off-the-shelf Stable Diffusion model [31] inpainting masked regions in real images – it correctly predicts shadows and supporting structures.
@@
To answer this question, we propose a general and lightweight evaluation protocol to systematically and efficiently ‘probe’ a pre-trained network on its ability to represent a number of ‘properties’ of the 3D scene and viewpoint. The protocol could be used for any network and any property of interest. We have probed properties including: 3D structure and material of the scene, such as surface layout; lighting, such as object-shadow relationships; and viewpoint dependent relations such as occlusion and depth.
@@
The protocol involves three steps: First, a suitable real image evaluation dataset is selected that contains ground truth annotations for the property of interest, for example the SOBA dataset [41] is used to probe the understanding of lighting, as it has annotations for object-shadow associations. This provides a train/val/test set for that property; Second, a grid search is carried out over the layers and time steps of the Stable Diffusion model, and transformer layers for other models, to select the optimal feature for determining that property. The selection involves learning the weights of a simple linear classifier for that property (e.g. ‘are these two regions in an object-shadow relationship or not’); Third, the selected feature (layer, time step) and trained classifier are evaluated on a test set, and its performance answers the question of how well the model ‘understands’ that property.
@@
In short, we probe scene geometry, material, support relation, shadow, occlusion and depth, to answer the question “To what extent do large vision models encode 3D properties of the scene?". We apply this protocol to a wide range of networks that are trained at large scale, including OpenCLIP [30, 18], DINOv1 [5], DINOv2 [27], and VQGAN [11]. This covers networks trained generatively (Stable Diffusion), with self-supervision (DINOv1 & DINOv2), with weak supervision (OpenCLIP), and by auto-regression (VQGAN).
@@
From our investigation, we make three observations: First, the Stable Diffusion and DINOv2 networks have a good ‘understanding’ of the scene geometry, support relations, the lighting, and the depth of a scene, with Stable Diffusion and DINOv2 having a similar and high prediction performance for these properties. However, their prediction of material and occlusion is poorer. Second, Stable Diffusion and DINOv2 generally demonstrate better performance for 3D properties than other networks trained at large scale: OpenCLIP, DINOv1, and VQGAN. Third, different time steps of Stable Diffusion features, as well as different transformer layers of DINO/CLIP/VQGAN, perform best for different 3D physical properties.
@@
Why is an understanding of the networks’ ability to predict 3D properties useful? There are four reasons: (1) It begins to answer the scientific question of the extent to which these networks implicitly model the 3D scene; (2) The features we determined that are able to predict 3D physical properties can be used for this task, e.g. to predict shadow-object associations or support relations. This could either be carried out directly by incorporating them in a prediction network, in the manner of [49]; or they can be used indirectly as a means to train a feed forward network to predict the properties [42, 43]; (3) By knowing what properties Stable Diffusion is not good at, we have a way to spot images generated by Stable Diffusion, as has been done by [32]; (4) It also reveals which properties the network could be trained further on to improve its 3D modelling, e.g., via extra supervision for that property.
2 Related Work
2.1 Exploration of Pre-trained Models
Building on the success of large-scale vision networks, there has been significant interest from the community to understand what has been learned by these complex models. On discriminative models, for example, [47, 26] propose inverse reconstruction to directly visualize the acquired semantic information in various layers of a trained classification network; [54, 13, 12] demonstrate that scene classification networks have remarkable localization ability despite being trained on only image-level labels; and [10, 37, 34] use the gradients of any target concept, flowing into the final convolutional layer to produce a saliency map highlighting important regions in the image for predicting the concept. In the more recent literature, [7] explores what has been learned in the powerful transformer model by visualizing the attention map. On generative models, researchers have mainly investigated what has been learned in GANs, for example, GAN dissection [3] presents an analytic framework to visualize and understand GANs at the unit-, object-, and scene-level; [44] analyse the latent style space of StyleGANs [20]. The most recent work [32] studies the 3D geometric relations in generated images, such as vanishing points and shadows, and notes that the errors made can be used to discriminate real from generated images. There is concurrent work [1] exploring the capability of predicting depth, surface normal and geometric correspondence for visual foundation models. In contrast to their work, we have studied a wider range of properties, covering both 3D geometric properties and 3D physical properties. Additionally, we have proposed a simple, general, yet efficient protocol for any property and any model and have investigated the performance of different time steps and layers of models for different properties.
2.2 Exploitation of Pre-trained Models
Apart from understanding the representation in pre-trained models, there has been a recent trend for exploiting models trained at large scale, to tackle a series of downstream tasks. For example, leveraging generative models for data augmentation in recognition tasks [19, 16], utilising large vision models for semantic segmentation [2, 46], open-vocabulary segmentation [23], depth map estimation [35, 53, 21, 50, 28], correspondence estimation [27, 51, 38, 25, 17] and pose estimation [52, 15]. More recently, [4] searched for intrinsic offsets in a pre-trained StyleGAN for a range of downstream tasks, predicting normal maps, depth maps, segmentations, albedo maps, and shading. In contrast to these works, we focus on exploring 3D physical understanding built into large models trained without 3D supervision.
2.3 3D Physical Scene Understanding
There have been works studying different 3D physical properties for scene understanding, including shadows [41, 40], material [39], occlusion [48], scene geometry [24], support relations [36] and depth [36]. However, these works focus on one or two physical properties, and most of them require training a model for the property in a supervised manner. In contrast, we use a single model to predict multiple properties, and do not train the features.
3 Method – Properties, Datasets, and Classifiers
Our goal is to examine the ability of large-scale vision models to understand different physical and geometrical properties of the 3D scene, including: scene geometry, material, support relations, shadows, occlusion and depth. Specifically, we conduct linear probing of the features from different layers and time steps of the Stable Diffusion model, and different transformer layers for other models including OpenCLIP, DINOv1, DINOv2 and VQGAN. First, we set up the questions for each property (Section 3.1); and then select real image datasets with ground truth annotations for each property (Section 3.2). We describe how a classifier is trained to answer the questions, and the grid search for the optimal time step and layer to extract a feature for predicting the property in Section 3.3.
3.1 Properties and Questions
Here, we study the large vision model’s ability to predict different properties of the 3D scene; the properties cover the 3D structure and material, the lighting, and the viewpoint. For each property, we propose questions that classify the relationship between a pair of Regions, A and B, in the same image, based on the features extracted from the large vision model. The properties and questions are:
  • Same Plane: ‘Are Region A and Region B on the same plane?’
  • Material: ‘Are Region A and Region B made of the same material?’
  • Shadow: ‘Are Region A and Region B in an object-shadow relationship?’
  • Depth: ‘Does Region A have a greater average depth than Region B?’
@@ 3.2 Datasets
Figure 2: Example images for probing scene geometry. The first row shows a sample annotation for the same plane, and the second row is a sample annotation for perpendicular plane. Here, and in the following figures, (A, B) are a positive pair, while (A, C) are negative. The images are from the ScanNetv2 dataset [8] with annotations for planes from [24]. In the first row, the first piece of floor (A) is on the same plane as the second piece of floor (B), but is not on the same plane as the surface of the drawers (C). In the second row, the table top (A) is perpendicular to the wall (B), but is not perpendicular to the stool top (C).
@@
Same Plane. We use the ScanNetv2 dataset [8] with annotations of planes from [24]. Regions are obtained via splitting plane masks into several regions. A pair of regions are positive if they are on the same plane, and negative if they are on different planes. The first row of Figure 2 is an example.
@@
Perpendicular Plane. We use the ScanNetv2 dataset [8]. We use the annotations from [24] which provide segmentation masks as well as plane parameters for planes in the image, so that we can obtain the normal of planes to judge whether they are perpendicular or not. A pair of regions are positive if they are on perpendicular planes, and negative if they are not on perpendicular planes. The second row of Figure 2 is an example.
Figure 3: Example images for probing material, support relation and shadow. The first row is for material, the second row for support relation, and the third row for shadow. First row: the material images are from the DMS dataset [39]. The paintings are both covered with glass (A and B) whereas the curtain (C) is made of fabric. Second row: the support relation images are from the NYUv2 dataset [36]. The paper (A) is supported by the table (B), but it is not supported by the chair (C). Third row: the shadow images are from the SOBA dataset [41]. The person (A) has the shadow (B), not the shadow (C).
@@
Material. We adopt the recent DMS dataset [39] to study the material property, which provides dense annotations of material category for each pixel in the images. Therefore, we can get the mask of each material via grouping pixels with the same material label together. In total, there are 46 pre-defined material categories. Regions are obtained by splitting the mask of each material into different connected components, i.e., they are simply groups with the same material labels, yet not connected. A pair of regions are positive if they are of the same material category, and negative if they are of different material categories. First row of Figure 3 is an example.
@@
Support Relation. We use the NYUv2 dataset [36] to probe the support relation. Segmentation annotations for different regions (objects or surfaces) are provided, as well as their support relations. Support relation here means an object is physically supported by another object, i.e., the second object will undertake the force to enable the first object to stay at its position. Regions are directly obtained via the segmentation annotations. A pair of regions are positive if the first region is supported by the second region, and negative if the first region is not supported by the second region. Second row of Figure 3 is an example.
@@
Shadow. We use the SOBA dataset [41, 40] to study the shadows which depend on the lighting of the scene. Segmentation masks for each object and shadow, as well as their associations are provided in the dataset annotations. Regions are directly obtained from the annotated object and shadow masks. In a region pair, there is one object mask and one shadow mask. A pair of regions are positive if the shadow mask is the shadow of the object, and negative if the shadow mask is the shadow of another object. Third row of Figure 3 is an example.
@@
Figure 4: Example images for probing viewpoint-dependent properties (occlusion & depth). The first row is for occlusion and the second row is for depth. First row: the occlusion images are from the Separated COCO dataset [48]. The sofa (A) and the sofa (B) are part of the same object, whilst the monitor (C) is not part of the sofa. Second row: the depth images are from the NYUv2 dataset [36]. The chair (A) is farther away than the object on the floor (B), but it is closer than the cupboard (C).
@@
Occlusion. We use the Seperated COCO dataset [48] to study the occlusion (object seperation) problem. Regions are different connected components of objects (and the object mask if it is not separated), i.e., groups of connected pixels belonging to the same object. A pair of regions are positive if they are different components of the same object separated due to occlusion, and negative if they are not from the same object. First row of Figure 4 is an example.
@@
Depth. We use the NYUv2 dataset [36], which provides mask annotations for different objects and regions, together with depth for each pixel. A pair of regions are positive if the first region has a greater average depth than the second region, and negative if the first region has a less average depth than the second region. The average depth of a region is calculated via the average of depth value of each pixel the region contains. Second row of Figure 4 is an example.